2x+7-5(x+1)=-(-4x-7)
Answer by user_dude2008(1861)

2x+7-5(x+1)=-(-4x-7)


2x+7-5x-5=4x+7

-3x+2=4x+7

-3x-4x=7-2

-7x=5

x=-5/7